linear algebra

Find the matrix A for T relative to the basis B. T: R2 R2, T(x, y) = (-3x + y, 3x - y), B = {(1, -1), (-1,5)} A =
Answer #1

Solution: linear operater T:R² R² such that Given Tlaly)= (-3nty, 3ary) T(1,-1)= (-3x101,311+1) = (-37, 3+1) = (-4,4) T(-1,5)(8-8) = a (1,-1) + ba (-1,5) = (92,-02)+(-b2, 5b2) (8,-8)= (az-br, 9, 15 br) ab-b2=80 -12 45 b2 =-8 - 6 462=0=162=0 from eall